# ChimeraLoRA: Multi-Head LoRA-Guided Synthetic Datasets

### v1

@@ Abstract
Beyond general recognition tasks, specialized domains including privacy-constrained medical applications and fine-grained settings often encounter data scarcity, especially for tail classes. To obtain less biased and more reliable models under such scarcity, practitioners leverage diffusion models to supplement underrepresented regions of real data. Specifically, recent studies fine-tune pretrained diffusion models with LoRA on few-shot real sets to synthesize additional images. While an image-wise LoRA trained on a single image captures fine-grained details yet offers limited diversity, a class-wise LoRA trained over all shots produces diverse images as it encodes class priors yet tends to overlook fine details. To combine both benefits, we separate the adapter into a class-shared LoRA $A$ for class priors and per-image LoRAs $\mathcal{B}$ for image-specific characteristics. To expose coherent class semantics in the shared LoRA $A$, we propose a semantic boosting by preserving class bounding boxes during training. For generation, we compose $A$ with a mixture of $\mathcal{B}$ using coefficients drawn from a Dirichlet distribution. Across diverse datasets, our synthesized images are both diverse and detail-rich while closely aligning with the few-shot real distribution, yielding robust gains in downstream classification accuracy.
@@ 1 Introduction
While general recognition tasks enjoy abundant and class-balanced data, specialized domains often face data scarcity [46, 25, 3] and long-tailed class distributions [27, 5]. For example, privacy constraints in medical domains and rarity in fine-grained tasks limit data collection, leaving only a few labeled images per class [35, 12]. Training under such data scarcity often causes models to overfit and learn decision boundaries biased toward majority classes, degrading generalization performance [2]. To supplement limited data, recent work leverages generative priors in pretrained text-to-image diffusion models [38] to synthesize additional training images by conditioning class names with text prompts [15]. However, without guidance from real images, synthetic data easily drift from the target distribution and lower downstream accuracy [44, 15].
To narrow the real-to-synthetic gap, recent work exploits few-shot real images [21, 20, 6, 15]. Specifically, a training-free image-wise line initializes the diffusion process from features of a single reference image to synthesize samples close to that reference [15]. Beyond this, an image-wise variant embeds the reference into a diffusion model by fine-tuning lightweight low-rank adapters (LoRAs) [17], enabling the model to capture fine-grained details [20]. However, these image-wise approaches make it difficult to generate diverse images, as they rely on a single image. To leverage the remaining images of the same class, a class-wise LoRA is fine-tuned on all shots to encode class-level priors and promote diversity, yet the resulting samples often overlook instance-specific details [21]. This trade-off stems from adapting the diffusion model with LoRA at a single granularity, either an image or a class, motivating a unified image- and class-level adaptation.
@@
  • Our synthetic datasets generally improve downstream accuracy across various benchmarks, including medical domains and long-tailed settings (Section 5.2).
@@ 3.2 Single-head LoRA-Guided Synthetic Datasets
Figure 3: Robust generation of semantic boosting (SB). Without SB, a LoRA trained on a one-shot image often fails to render a car even when prompted with “a photo of a car”. With SB, repeated exposure to the car region during training robustly generates complete cars.
@@ 4.2 LoRA Merging for Image Generation
Here, the concentration $\alpha$ controls how the mixture spreads over the simplex $\Delta^{K-1}$. When $\alpha=1$, $\mathbf{w}$ is uniformly distributed over the simplex. For $\alpha<1$, the distribution becomes sparse and typically concentrates most of its mass on a single $B_{i}$, which behaves like an image-wise regime. When $\alpha>1$, the weights cluster near the uniform vector and approximate a class-wise regime. We refer to this framework as ChimeraLoRA, which establishes a class-level backbone from the few-shot references and attaches instance-specific details to produce coherent semantics, yet diverse images.
@@
When $\alpha=1$, $\mathbf{w}$ is uniform on the simplex $\Delta^{K-1}$, and we observe that this setting typically yields decent downstream performance. However, sampling a fresh $\mathbf{w}$ for every image can hinder batch-wise generation, as $B^{\prime}$ must be rebuilt per sample. We consider two practical variants to mitigate this overhead: (i) set $w_{i}=1/K$ in (5), which still maintains high fidelity with reasonable coverage, as shown in Figure 2; and (ii) reuse a single $\mathbf{w}\sim\text{Dirichlet}(\mathbf{1})$ to synthesize multiple images, with diffusion stochasticity providing additional variation. However, we note that per-image mixtures with $\text{Dirichlet}(\mathbf{1})$ still provide the broadest coverage.
@@ 5.1 Experimental Setup
We evaluate on 11 publicly available image classification datasets: FGVCAircraft (AIR) [30], Caltech101 (CAL) [9], StanfordCars (CAR) [23], DTD [4], EuroSAT (EUR) [16], Flowers102 (FLO) [32], Food101 (FOD) [1], OxfordPets (PET) [33], Skin Lesions (ISIC) [19], CIFAR-10 [24], and ImageNet100 [40]. Our benchmarks cover diverse fine-grained tasks including cars and pets, and also specialized domains such as satellite imagery, textures, and medical dermatology, reflecting practical few-shot constraints encountered in real applications. Instead of using the full training split, we randomly sample four real images per class as references and evaluate on the standard test splits.
Table 1: Downstream performance with synthetic datasets under few-shot scenarios. Starting with 4-shot labels, we generate 500 additional images per class to train on 504 per class in total, improving accuracy by 2.1 percentage points on average over state-of-the-art baselines across nine datasets. We mark the best in bold and the second best with underlines.
@@
Long Tail Avg. Long Tail Avg. Long Tail Avg. Long Tail Avg. Long Tail Avg.
@@
We compare our method against three methods: IsSynth [14], LoFT [20], and DataDream [21]. IsSynth is train-free and synthesizes data using features extracted from the given 4-shot references. LoFT and DataDream fine-tune diffusion models with LoRA in image-wise and class-wise configurations, respectively. To match trainable-parameter budgets in the 4-shot setting, DataDream uses LoRA rank 16, since it trains a single class-wise LoRA over all four images, whereas LoFT and our method use rank 4, as the adaption is split across the four images. We note that thanks to the shared adapter $A$, our approach uses 37.5% fewer trainable parameters than both baselines. We employ a multi-head LoRA with distinct learning rates: $1\times 10^{-4}$ for LoRA $A$ and $1\times 10^{-3}$ for LoRA $B$. All LoRA-based methods use Stable Diffusion 2.1 [38] as the base diffusion model. For generation, the trained adapters are attached to Stable Diffusion with guidance scale $2$, and when composing the per-image adapters we draw mixture coefficients from $\text{Dirichlet}(\mathbf{1})$.
@@ 6 Conclusion
When the few-shot references share similar semantics, our method can synthesize images aligned with the few-shot reference distribution and improve downstream tasks, but it is difficult to apply when the references differ semantically. With few-shot images that span different semantics, the framework can be extended to generate per-semantic samples as well as intermediate or more diverse variants. These can serve as soft-label augmentations, such as images with soft labels, and are expected to further benefit downstream tasks.

### v2

@@ Abstract
Beyond general recognition tasks, specialized domains and fine-grained settings often encounter data scarcity, especially for tail classes. To obtain less biased and more reliable models under such scarcity, practitioners leverage diffusion models to supplement underrepresented regions of real data. Specifically, recent studies fine-tune pretrained diffusion models with LoRA on few-shot real sets to synthesize additional images. While an image-wise LoRA trained on a single image captures fine-grained details yet offers limited diversity, a class-wise LoRA trained over all shots produces diverse images as it encodes class priors yet tends to overlook fine details. To combine both benefits, we separate the adapter into a class-shared LoRA $A$ for class priors and per-image LoRAs $\mathcal{B}$ for image-specific characteristics. To expose coherent class semantics in the shared LoRA $A$, we propose a semantic boosting by preserving class bounding boxes during training. For generation, we compose $A$ with a mixture of $\mathcal{B}$ using coefficients drawn from a Dirichlet distribution. Across diverse datasets, our synthesized images are both diverse and detail-rich while closely aligning with the few-shot real distribution, yielding robust gains in downstream classification accuracy.
@@ 1 Introduction
While general recognition tasks enjoy abundant and class-balanced data, specialized domains often face data scarcity [46, 25, 3] and long-tailed class distributions [27, 5]. For example, class rarity in fine-grained tasks can limit data collection, leaving only a few labeled images per class [35, 12]. Training under such data scarcity often causes models to overfit and learn decision boundaries biased toward majority classes, degrading generalization performance [2]. To supplement limited data, recent work leverages generative priors in pretrained text-to-image diffusion models [38] to synthesize additional training images by conditioning class names with text prompts [15]. However, without guidance from real images, synthetic data easily drifts from the target distribution and lower downstream accuracy [44, 15].
To narrow the real-to-synthetic gap, recent work exploits few-shot real images [21, 20, 6, 15]. Specifically, a training-free image-wise baseline initializes the diffusion process from features of a single reference image to synthesize samples close to that reference [15]. Beyond this, an image-wise variant embeds the reference into a diffusion model by fine-tuning lightweight low-rank adapters (LoRAs) [17], enabling the model to capture fine-grained details [20]. However, these image-wise approaches make it difficult to generate diverse images, as they rely on a single image. To leverage the remaining images of the same class, a class-wise LoRA is fine-tuned on all shots to encode class-level priors and promote diversity, yet the resulting samples often overlook instance-specific details [21]. This trade-off stems from adapting the diffusion model with LoRA at a single granularity, either an image or a class, motivating a unified image- and class-level adaptation.
@@
  • Our synthetic datasets generally improve downstream accuracy across various benchmarks, including specialized domains and long-tailed settings (Section 5.2).
@@ 3.2 Single-head LoRA-Guided Synthetic Datasets
Figure 3: Robust generation using semantic boosting (SB). Without SB, a LoRA trained on a one-shot image often fails to render a car even when prompted with “a photo of a car”. With SB, repeated exposure to the car region during training robustly generates complete cars.
@@ 4.2 LoRA Merging for Image Generation
Here, the concentration $\alpha$ controls how the mixture spreads over the simplex $\Delta^{K-1}$. When $\alpha=1$, $\mathbf{w}$ is uniformly distributed over the simplex. For $\alpha<1$, the distribution becomes sparse and typically concentrates most of its mass on a single $B_{i}$, which behaves like an image-wise regime. When $\alpha>1$, the weights cluster near the uniform vector and approximate a class-wise regime. We refer to this framework as ChimeraLoRA, which establishes a class-level backbone from the few-shot references and attaches instance-specific details to produce coherent semantics, yet diverse images. [^1] In the Appendix A, we expand our discussion to the setting with two concentration parameters, $\alpha$ and $\beta$, rather than a single $\alpha$.
@@
When $\alpha=1$, $\mathbf{w}$ is uniform on the simplex $\Delta^{K-1}$, and we observe that this setting typically yields decent downstream performance. However, sampling a fresh $\mathbf{w}$ for every image can hinder batch-wise generation, as $B^{\prime}$ must be rebuilt per sample. We consider two practical variants to mitigate this overhead: (i) set $w_{i}=1/K$ in (5), which still maintains high fidelity with reasonable coverage, as shown in Figure 2; and (ii) reuse a single $\mathbf{w}\sim\text{Dirichlet}(\mathbf{1})$ to synthesize multiple images, with diffusion stochasticity providing additional variation. However, we note that per-image mixtures with $\text{Dirichlet}(\mathbf{1})$ still provide the broadest coverage. In Appendix B.2, we analyze the trade-off between wall-time and accuracy across the three methods.
@@ 5.1 Experimental Setup
We evaluate on 11 publicly available image classification datasets: FGVCAircraft (AIR) [30], Caltech101 (CAL) [9], StanfordCars (CAR) [23], DTD [4], EuroSAT (EUR) [16], Flowers102 (FLO) [32], Food101 (FOD) [1], OxfordPets (PET) [33], Skin Lesions (ISIC) [19], CIFAR-10 [24], and ImageNet100 [40]. Our benchmarks cover diverse fine-grained tasks including cars and pets, and also specialized domains such as satellite imagery, textures, and medical dermatology, reflecting practical few-shot constraints in real applications.
Table 1: Downstream performance with synthetic datasets under 4-shot scenarios. Starting with 4-shot labels, we generate 500 additional images per class to train on 504 per class in total, improving accuracy by 2.1 percentage points on average over state-of-the-art baselines across nine datasets. We mark the best in bold and the second best with underlines.
@@
Head Tail Avg. Head Tail Avg. Head Tail Avg. Head Tail Avg. Head Tail Avg.
@@
We compare our method against three methods: IsSynth [14], LoFT [20], and DataDream [21]. IsSynth is train-free and synthesizes data using features extracted from the given 4-shot references. LoFT and DataDream fine-tune diffusion models with LoRA in image-wise and class-wise configurations, respectively. To match trainable-parameter budgets in the 4-shot setting, DataDream uses LoRA rank 16, since it trains a single class-wise LoRA over all four images, whereas LoFT and our method use rank 4, as the adaption is split across the four images. We note that thanks to the shared adapter $A$, our approach uses 37.5% fewer trainable parameters than both baselines. We employ a multi-head LoRA with distinct learning rates: $1\times 10^{-4}$ for LoRA $A$ and $1\times 10^{-3}$ for LoRA $B$. All LoRA-based methods use Stable Diffusion 2.1 [38] as the base diffusion model. For generation, the trained adapters are attached to Stable Diffusion with guidance scale $2$, and when composing the per-image adapters we draw mixture coefficients from $\text{Dirichlet}(\mathbf{1})$. Unless otherwise noted, we reproduce the results for all baselines.
@@ 5.4 Further Ablation Studies
We conduct a range of ablation studies on our method, and additional ablation results are provided in the Appendix C.
@@ 6 Conclusion
We use Grounded-SAM as a general solution across domains for Semantic Boosting, but in medical settings, domain-specific tools such as MedSAM may be more appropriate, and our current medical domain validation remains limited. Future work will strengthen domain-specific evidence by evaluating on additional medical datasets, conducting robustness analyses under clinically relevant perturbations, and expanding the discussion of related work on generative augmentation for long-tailed medical image classification rather than assuming that simple tool substitution is sufficient.
Acknowledgements.
This work was partly supported by the IITP grants and the NRF grants funded by Ministry of Science and ICT, Korea (No.RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); No.RS-2024-00457882, AI Research Hub Project; IITP-2026-RS-2024-00437866; RS-2024-00509258, Global AI Frontier Lab).
Supplementary Material
A Experiments on 16-Shot Setting
Table LABEL:app:main-table shows that our method works well even in the 16-shot setting. Rather than using a single parameter $\alpha$, we employ two concentration parameters, $\alpha$ and $\beta$, to define the Dirichlet distribution. Specifically, we sample weights $(w_{1},\dots,w_{16})\sim\mathrm{Dir}(\mathbf{c})$, where $\mathbf{c}\in\mathbb{R}^{16}$ is defined as:
$c_{i}=\begin{cases}\alpha&\text{if }i=k\\ \beta&\text{if }i\neq k\end{cases},\quad k\sim\text{Uniform}(\{1,\dots,16\})\;.$ (8)
We set $\alpha=2$ and $\beta=0.07$, yielding $\mathbb{E}[w_{k}]=\frac{\alpha}{\alpha+(16-1)\beta}\approx 66\%$ for the main adapter, with the remaining $\approx 34\%$ mass distributed over the other adapters. Choosing a small $\beta<1$ promotes sparsity among the remaining adapters, i.e., the residual mass is typically concentrated on only a few adapters rather than being spread uniformly across all of them. In summary, as the number of shots increases, using a relatively larger $\alpha$ (relative to $\beta$) becomes beneficial for capturing fine-grained details by placing more emphasis on the main adapter, while still preventing too many adapters from being mixed in.
B Computation Analyses
B.1 Resource Usage
As each image has its own LoRA head, resource usage naturally grows as the number of shots increases, leading to higher memory consumption and more trainable parameters. To keep this overhead manageable, we reduce the LoRA rank for multi-shot settings. In particular, the results reported in Table 1 for the 4-shot setting are obtained by setting the rank to one quarter of the baseline, which substantially lowers the per-head footprint while retaining competitive performance. Table 6 further summarizes the resulting parameter counts and training time, showing that our approach offers a practical middle ground between baselines in terms of both memory and speed.
Method Rank # parameters Training time
DataDream 16 3.2M 95s
LoFT 4 0.8M 356s
Ours 4 1.9M 212s
Table 6: Trade-off between # parameters and training time.
B.2 Effect of Weight Sampling on Efficiency
In Section 4.2, when $\alpha=1$ we draw $\mathbf{w}\sim\mathrm{Dirichlet}(\mathbf{1})$, i.e., $\mathbf{w}$ is uniformly distributed over the simplex $\Delta^{K-1}$, and we compare three practical choices: Uniform $w$ (fix $w_{i}=1/K$), Reuse $w$ (reuse a single $\mathbf{w}\sim\mathrm{Dirichlet}(\mathbf{1})$ for multiple images), and New $w$ (sample a fresh $\mathbf{w}$ per image). Table 7 shows that moving from Uniform $\rightarrow$ Reuse $\rightarrow$ New increases the diversity of per-image mixtures and tends to improve accuracy, but also increases wall-time, where wall-time denotes the overall generation time, since per-image sampling reduces batching efficiency.
Method Wall-Time Accuracy
Uniform $w$ 4.6h 59.37
Reuse $w$ 5.8h 61.31
New $w$ 6.1h 61.51
Table 7: Trade-off between wall-time and accuracy.
C Additional Ablation Studies
C.1 Rank $r$ Ablation
On the DTD dataset, we observe that increasing the LoRA rank $r$ yields modest but consistent gains: with the number of synthetic images fixed at 100, performance improves from 55.07 at r=2r{=}2 to 57.20 and 57.72 at r=8r{=}8 and r=16r{=}16, respectively. This trend aligns with the increased capacity of higher-rank adapters, but it also comes with higher computational cost, as larger $r$ increases the number of trainable parameters and the associated memory and training-time overhead. Overall, we find that higher ranks provide incremental improvements, reflecting a typical trade-off between computation resources and performance.
C.2 $\alpha$ Ablation
In a symmetric Dirichlet distribution, $\alpha$ primarily controls the concentration of the mixture weights $\mathbf{w}$: smaller $\alpha$ produces spikier weights that are closer to selecting a single head, whereas larger $\alpha$ spreads mass more evenly and behaves like a more uniform ensemble. As summarized in Table 8, this induces a clear trade-off. With smaller $\alpha$, synthesis tends to be more faithful and yields better FID, while larger $\alpha$ encourages broader mixing across heads and improves diversity, reflected by a higher distance score. Overall, $\alpha$ serves as a simple tunable parameter to balance fidelity and diversity depending on the desired operating point.
Metric $\alpha=0.2$ $\alpha=1.0$ $\alpha=5.0$
FID (Real vs. Synthetic, $\downarrow$) 0.098 0.101 0.146
Distance (Inter-class, $\uparrow$) 0.142 0.153 0.199
Table 8: Comparison metrics under different $\alpha$ values.
C.3 Semantic Boosting on Baselines
The table 9 summarizes the effect of applying semantic boosting to the baselines. Overall, semantic boosting consistently improves performance, and the gain is most pronounced for image-wise LoRA training, where optimization can otherwise overfit to spurious regions or drift away from the target instance.
Method AIR FLO
DataDream (+SB) 44.3 $\rightarrow$ 44.2 92.9 $\rightarrow$ 93.1
LoFT (+SB) 41.7 $\rightarrow$ 44.4 91.3 $\rightarrow$ 92.2
Ours 43.9 $\rightarrow$ 46.0 93.1 $\rightarrow$ 93.4
Table 9: Effect of semantic boosting on baselines.
D Additional Qualitative Results
Figures 9 and 10 present additional qualitative results across diverse datasets.
Figure 9: Additional Qualitative Results. (a) LoFT, (b) DataDream, (c) ChimeraLoRA. From top to bottom, the classes are 1991 Volkswagen Golf Hatchback and 2007 BMW X5 SUV from Cars, followed by annualcrop and highway from EuroSAT.
Figure 10: Additional Qualitative Results. (a) LoFT, (b) DataDream, (c) ChimeraLoRA. From top to bottom, the classes are banded and bubbly from DTD, followed by basal cell carcinoma and vascular lesion from ISIC.